# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v1

@@ 2.2.2 Low-rank branch compensation
Symmetric quantization forces the zero point to align with the floating point, resulting in a quantization range that is symmetric with respect to zero. The quantization process is typically defined as:
$X_{q}=\text{round}\left(\frac{X}{\Delta}\right),\quad\Delta=\frac{\max(|X|)}{2^{N}-1}$ (12)
where the floating-point tensor $X$ is scaled by a parameter $\Delta$, which is computed from the distribution of $X$. Eq. (12) presents a commonly used Min-Max method for computing the quantization step size $\Delta$, where $N$ denotes the bit-width of the quantization precision.
The dequantization process can be formulated as:
$\text{dequant}(X)=X_{q}\Delta$ (13)
$\text{dequant}(Y_{q})=X_{q}W_{q}\Delta_{X}\Delta_{W}=Y_{q}\Delta_{X}\Delta_{W}$ (14)
It is evident that symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output with the product of the scale factors for the weights and activations like Eq. (14). This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization allows the zero point to shift dynamically according to the minimum and maximum values of the data. This enables a tighter fit to the data distribution, especially when it is skewed, at the cost of slightly more complex computation.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy. The dequantization process can be formulated as:
$\text{dequant}(X)=(X_{q}-Z)\Delta$ (15)
$\text{dequant}(Y_{q})=(X_{q}-Z_{X})\Delta_{X}\cdot(W_{q}-Z_{W})\Delta_{W}$
$=Y_{q}\Delta_{X}\Delta_{W}+(Z_{X}W_{q}+Z_{W}X_{q}+Z_{X}Z_{W})\Delta_{X}\Delta_{W}$ (16)
As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
2.4 Granularity of quantization
The granularity of quantization often largely effects the low-bit quantization performance. Commonly adopted granularities include per-tensor, per-channel or per-token, and per-group, as illustrated in Fig. 4.
@@ 2.4 Granularity of quantization
In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor minimum and maximum values. A finer-grained alternative is per-channel quantization for weights and per-token quantization for activations, where each input channel of the weights or each token of the activations shares the same quantization parameters. For low-bit quantization scenarios, even finer granularity, per-group quantization, is often employed to further enhance performance. In this case, the elements within a single channel or token are subdivided into groups of a predefined size, with each group quantized independently using its own parameters.
Finer granularity typically leads to improved performance but also incurs higher memory overhead due to the need to store more quantization parameters. Besides, as QServe [17] mentioned, per-group quantization will also raise computational cost in hardware level. Therefore, while finer granularity can enhance accuracy, it may also reduce the overall efficiency of the system. In weight-only quantization, papers such as GPTQ [8], AWQ [11], and QuIP [9] all employ per-group quantization to further enhance performance. For weight-activation quantization, Atom [1] also adopts a per-group quantization strategy. However, with the introduction of rotation-based methods in QuaRot [13] and SpinQuant [3], the performance of per-channel quantization has been significantly improved. Inspired by this, subsequent papers like FlatQuant [12], ResQ [19], and OSTQuant [4] have also adopted per-channel granularity for quantization.
2.5 MXFP4 data format
Recently, the NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [22] is released, which has enabled very low-precision operations using FP4 variants like MXFP4 [23, 24] and NVFP4 [25]. MXFP4 stores each value in an E2M1 format, consisting of one sign bit, one mantissa bit, and two exponent bits. Unlike the INT4 format, MXFP4 employs a fixed granularity, utilizing symmetric per-group quantization with a group size of 32. Each group is associated with an E8M0 scaling factor, and this rescaling process is hardware-integrated.
The value of an encoding MXFP4 data is represented as:
@@ 2.5 MXFP4 data format
where $S,E,M$ corresponds to the sign,exponent and mantissa fields, respectively. The bias is the exponent bias which is set to 1 for MXFP4. So the MXFP4 encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
The quantization process of MXFP4 is similar to the integer quantization process. First, we compute the scaling factor, using the same way as symmetric quantization situation mentioned above that computes from the maximum value of a group of values. Then the scaling factors need to be cast to E8M0 format. After scaling, the values are quantized to the nearest MXFP4 encoding values.
MXFP4 is enable to better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, and MXFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, MXFP4 is more suitable than INT4 as a quantization data type.
@@ 3 Extensive Evaluations
Quantization settings We conduct experiments using W4A4 precision, comparing the results with the origin precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
@@
Implementation. To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank branch, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
@@ 3.1 Overall performance
DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 3: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
@@
DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 4: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
We select two of the most widely-used pre-quantization transformation techniques, scaling and rotation, and combined them with two quantization error mitigation methods: GPTQ and low-rank branch compensation. A series of experiments are conducted to evaluate these combinations. Table 3 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations. Table 4 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank branch. This configuration yields a perplexity of 11.73 in per-channel quantization for Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
@@
In addition, using both optimized rotation and scaling improves performance over using only optimized rotation. However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not performs well together with rotation in W4A4 settings.
Model precision Wiki2 PPL($\downarrow$) Zero-Shot Accuracy($\uparrow$)
@@
Table 5: Comparision of s ymmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
In Table 5, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank branch.
@@
Table 6: Comparision of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [32] accuracy of LlaMA-3.2-1B
@@
Therefore, an effective configuration is to use symmetric quantization for weights and asymmetric quantization for activations. This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with dequantization process, compared to fully asymmetric quantization for both weights and activations.
@@ 3.3 Granularity
As illustrated in Fig. 6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the FP16-format scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 7, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 6: Comparision of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
@@
Average Extra bits 0.5 0.25 0.125 0.0625 0.03125
@@
Table 7: Average extra bits introduced by scaling factors of different granularities
3.4 MXFP4 vs. INT4
Extending experiments to the MXFP4 format, we follow MXFP4 configurations implementing symmetric quantization with group size 32. For activations, we use a clipping threshold of 3/4. For weights, we follow the experiments of INT4 quantization, using a linear search over the squared error to find a best clipping ratio. First, we apply no extra pre-quantization transformation or quantization error mitigation process, conducting experiments using only RTN (round to nearest). Table 8 demonstrates that, MXFP4 shows much better performance than INT4 (per-channel).
Error compensation Data Format
INT4 (channel) INT4 (g-32) MXFP4
- 923.72 14.19 15.19
GPTQ 1007.33 14.73 13.18
low-rank 723.54 14.04 15.59
GPTQ+low-rank 578.61 12.48 12.72
Table 8: Evaluation of MXFP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
Evaluate Methods WikiText2PPL($\downarrow$)
Data Normalization Error Compensation INT4(channel) MXFP4
scaling GPTQ 134.85 13.06
scaling low-rank 165.4 14.83
scaling GPTQ+low-rank 124.56 12.64
optimized rotation+scaling GPTQ 11.8 12.61
optimized rotation+scaling low-rank 14.58 15.53
optimized rotation+scaling GPTQ+low-rank 11.73 12.29
Table 9: Evaluation of MXFP4 format quantization, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
Further, we conduct experiments using both pre-quantization transformation and quantization error mitigation techniques. Results in Table 9 demonstrate that both pre-quantization transformation and quantization error mitigation techniques achieve competitive performance under MXFP4 quantization, confirming the direct applicability of INT4 quantization methods to FP4. Crucially, scaling-based approaches proved more effective than rotation-based methods for MXFP4 –- a reversal of their relative efficacy observed in INT4 quantization scenarios.

### v2

@@ 2.2 Settings of quantization
Symmetric quantization forces the zero point to align with the floating point, resulting in a quantization range that is symmetric with respect to zero. Asymmetric quantization allows the zero point to shift dynamically according to the minimum and maximum values of the data. This enables a tighter fit to the data distribution, especially when it is skewed, at the cost of slightly more complex computation.
Symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy. As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
2.2.2 Granularity of quantization
The granularity of quantization often largely affects the low-bit quantization performance. Commonly adopted granularities include per-tensor, per-channel, or per-token, and per-group, as illustrated in Fig. 4.
@@ 2.2.2 Granularity of quantization
In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor’s minimum and maximum values. A finer-grained alternative is per-channel quantization for weights and per-token quantization for activations, where each input channel of the weights or each token of the activations shares the same quantization parameters. For low-bit quantization scenarios, even finer granularity, per-group quantization, is often employed to further enhance performance. In this case, the elements within a single channel or token are subdivided into groups of a predefined size, with each group quantized independently using its own parameters.
Finer granularity typically leads to improved performance but also incurs higher memory overhead due to the need to store more quantization parameters. Besides, as QServe [15] mentioned, per-group quantization will also raise computational cost at the hardware level. Therefore, while finer granularity can enhance accuracy, it may also reduce the overall efficiency of the system.
In weight-only quantization, papers such as GPTQ [8], AWQ [11], and QuIP [9] all employ per-group quantization to further enhance performance. For weight-activation quantization, Atom [1] also adopts a per-group quantization strategy. However, with the introduction of rotation-based methods in QuaRot [13] and SpinQuant [3], the performance of per-channel quantization has been significantly improved. Inspired by this, subsequent papers like FlatQuant [12], ResQ [17], and OSTQuant [4] have also adopted per-channel granularity for quantization.
2.3 FP4 Quantization and Variant Formats
Recently, NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [24] has been released, which has enabled very low-precision operations using FP4 variants like MXFP4 [25, 26] and NVFP4 [27]. Both MXFP4 and NVFP4 store each value in an E2M1 format, which consists of one sign bit, one mantissa bit, and two exponent bits. Unlike the INT4 format, they adopt a fixed granularity and use symmetric per-group quantization, with MXFP4 employing a group size of 32 and NVFP4 using a finer granularity of 16. Each group is associated with a scaling factor, and this rescaling process is hardware-accelerated. MXFP4 uses an FP8 (E8M0) format, while NVFP4 uses an FP8 (E4M3) format. Additionally, NVFP4 incorporates an extra FP32 scaling factor per tensor to extend its representable range.
The value of an encoding FP4 (E2M1) data is represented as:
@@ 2.3 FP4 Quantization and Variant Formats
where $S,E,M$ correspond to the sign, exponent, and mantissa fields, respectively. The bias is the exponent bias, which is set to 1. So the FP4 (E2M1) encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
FP4 can better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, and MXFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, FP4 is more suitable than INT4 as a quantization data type.
@@ 3 Extensive Evaluations
Quantization settings We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
@@
Implementation. To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows the OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank decomposition, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
@@ 3.1 Overall performance
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 8: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
@@
PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
@@
Table 9: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B. PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank decomposition compensation. A series of experiments is conducted to evaluate these combinations. Table 8 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations. Table 9 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank decomposition. This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
@@
In addition, using both optimized rotation and scaling improves performance over using only optimized rotation. However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not perform well together with rotation in W4A4 settings.
Model precision
Wiki2
PPL($\downarrow$)
Zero-Shot Accuracy($\uparrow$)
@@
Table 10: Comparison of symmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
In Table 10, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank decomposition.
@@
Table 11: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [34] accuracy of LlaMA-3.2-1B
@@
Therefore, an effective configuration is to use symmetric quantization for weights and asymmetric quantization for activations. This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with the dequantization process, compared to fully asymmetric quantization for both weights and activations.
@@ 3.3 Granularity
As illustrated in Fig. 6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 12, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 6: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
@@
Average Extra bits
1 0.5 0.25 0.125 0.0625
@@
Table 12: Average extra bits introduced by FP32 scaling factors of different granularities
3.4 FP4 vs. INT4
Extending our experiments to the FP4 format, we compare several variants of FP4, including MXFP4, NVFP4, and naive FP4 implementation with varying granularities and scaling factor formats. For per-channel quantization, we employ asymmetric quantization for activations and symmetric quantization for weights. To ensure a fair comparison with MXFP4 and NVFP4, all other experiments using per-group quantization adopt symmetric quantization for both activations and weights.
First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods. Table 13 demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel). We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results. For finer granularity scenarios, FP4 and INT4 show very small differences, and both the format of the scaling factor and the size of the group significantly impact performance. Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format. With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision. By contrast, MXFP4 uses the E8M0 format, which restricts scale factors to powers of two, resulting in much coarser quantization. Moreover, NVFP4 employs a two-level micro-block scaling strategy: it first applies a per-tensor scaling factor (FP32) to compensate for the smaller range that E4M3 scales can represent. NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [40].
Quantization
error
mitigation
Data Format
INT4
channel, FP32
FP4
channel, FP32
INT4
g16, FP32
FP4
g16, FP32
MXFP4
g32, E8M0
FP4
g32, E4M3
FP4
g16, E8M0
NVFP4
g16, E4M3
RTN 256.97 135.68 11.75 11.34 15.17 12.04 14.52 11.54
GPTQ 161.04 101.36 11.47 10.87 13.24 11.36 12.77 10.98
low-rank 179.87 132.14 11.69 11.32 15.56 11.98 14.67 11.48
GPTQ+low-rank 139.31 96.33 10.92 10.76 12.69 11.18 12.25 10.83
Table 13: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor. And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
Quantization
error
mitigation
Data Format
NVFP4
W:E4M3, A:E4M3
FP4
W:E2M1, A:E4M3
FP4
W:E2M1, A:E2M1
FP4
W:INT4, A:E4M3
FP4
W:INT4, A:INT4
RTN 11.54 14.10 4130 14.66 4777
GPTQ 10.98 13.38 3618 17.96 3803
Low-rank 11.48 13.75 4057 14.43 5356
GPTQ+Low-rank 10.83 11.38 3650 11.59 4124
Table 14: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. E4M3, E2M1, and INT4 refer to the format of the scaling factor. W and A denote weight and activation, respectively. All of these follow NVFP4’s approach, employing a two-level scaling scheme.
We further reduce the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting. As shown in Table 14, applying 4-bit scaling factors to both weights and activations leads to very poor performance. When applied only to weights, the performance remains moderate and is worse compared with NVFP4.
Evaluate Methods Wiki2‑PPL($\downarrow$)
Pre-quantization
transformation
Quantization error
mitigation
INT4
channel, FP32
MXFP4
g32, E8M0
NVFP4
g16, E4M3
scaling GPTQ 134.85 13.24 10.97
scaling low-rank 165.4 14.91 11.36
scaling GPTQ+low-rank 124.56 12.58 10.81
optimized rotation+scaling GPTQ 11.8 12.37 11.04
optimized rotation+scaling low-rank 14.58 15.43 12.57
optimized rotation+scaling GPTQ+low-rank 11.73 12.29 11.01
Table 15: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor.
Notably, rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel). As argued in AMXFP4 [41], grouping activation tensors into small micro-scaled units mitigates outliers like rotation. For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.